OpenStudy (anonymous):

help me:(((( 1. Solve 16x + 8 = 4y for y. Put your answer in the form y=mx+b with no spaces. y=-4x+2 y=4x+2 y=4x-2 y=-4x-2 2. Solve -28x = 4y + 16 for y. Put your answer in the form y=mx+b with no spaces. y=7x+4 y=-7x+4 y=7x-4 y=-7x-4 3. Solve 8y = 64x - 80 for y. Put your answer in the form y=mx+b with no spaces. y=8x-10 y=8x+10 y=-8x+10 y=-8x-10

OpenStudy (jim766):

solve for y, means to get it by y alone.... 16x + 8 = 4y what is the only thing keep the y from being by iteself?

OpenStudy (anonymous):

you just made me more confused

jimthompson5910 (jim_thompson5910):

I'll do the first one to get you started. The idea is to solve for y or isolate it completely. This means that you get the y term all alone on one side (and everything else is on the other side) 16x + 8 = 4y 4y = 16x + 8 y = 16x/4 + 8/4 ... divide EVERY term by 4 to isolate y y = 4x + 2

jimthompson5910 (jim_thompson5910):

So 16x + 8 = 4y turns into y = 4x + 2 which is now in y = mx + b form where m = 4 (slope) b = 2 (the y-intercept, which as a point is (0, 2))